OpenStudy (anonymous):

Find the point on the terminal side of θ = -3pi/4 that has an x coordinate of -1. Show your work for full credit

OpenStudy (anonymous):

\[\tan\theta=\frac{opp}{adj}\] \[opp = adj\times \tan\theta\] \[opp = 1\times \tan(\frac{\pi}{4})\] \[point = (-adj,-opp)\]